P. 248 Given: L (5² S² 2 Criven: L{ 52, Convolution theorem
Answer #1

Soli 서 cet fcs) - 4 8(s) - s L-1888 = L(a) = Cost = ft) L {8} = L ( t) - Cost = $it) using convolution theorem to Life) f(s)]Pape ।। sh (571)2 Sinat 2 ary